# LEKA: LLM-Enhanced Knowledge Augmentation

### v1

@@ Abstract
Humans excel in analogical learning and knowledge transfer and, more importantly, possess a unique understanding of identifying appropriate sources of knowledge. From a model’s perspective, this presents an interesting challenge. If models could autonomously retrieve knowledge useful for transfer or decision-making to solve problems, they would transition from passively acquiring to actively accessing and learning from knowledge. However, filling models with knowledge is relatively straightforward—it simply requires more training and accessible knowledge bases. The more complex task is teaching models about which knowledge can be analogized and transferred. Therefore, we design a knowledge augmentation method LEKA for knowledge transfer that actively searches for suitable knowledge sources that can enrich the target domain’s knowledge. This LEKA method extracts key information from textual information from the target domain, retrieves pertinent data from external data libraries, and harmonizes retrieved data with the target domain data in feature space and marginal probability measures. We validate the effectiveness of our approach through extensive experiments across various domains and demonstrate significant improvements over traditional methods in reducing computational costs, automating data alignment, and optimizing transfer learning outcomes.
Introduction
@@
Humans are good at identifying proper sources of knowledge. This is an ability that is rooted in our capability for analogical reasoning and knowledge management. On the contrary, artificial intelligence models do not inherently possess this intuition: they require explicit instructions and systematic training to identify and utilize relevant information. This gap presents a challenge in enriching domain knowledge and enhancing data augmentation.
@@
Knowledge augmentation is vital for improving model performance, especially in domains with limited information or complex data structures (Tang et al. 2020). In knowledge augmentation, knowledge transfer is an important method that improves learning performance by transferring knowledge from external information sources. In data-limited scenarios, effective knowledge sourcing can bridge domain information gaps and enhance model robustness by leveraging relevant information from external sources. In these scenarios, knowledge augmentation, especially knowledge transfer can reduce reliance on extensive target domain data by strategically selecting sources matching the target needs.
@@
Despite its potential, traditional knowledge augmentation methods often involve manual intervention, where human experts pick source domains based on their subjective interpretation of the target domain’s requirements. Using tabular learning as an example highlights several challenges: (1) structural and format differences across domains hinder data alignment and integration; (2) discrepancies in tasks and content between the selected source and target domains can diminish the success of knowledge augmentation; (3) extensive data preprocessing is often required to match the chosen source domain dataset properly. Additionally, relying on human expertise for source domain selection can lead to biases and inefficiencies, as these choices are typically made based on prior knowledge rather than through a rigorous, data-driven analysis.
@@
A natural idea is to construct an automated search for datasets with relevant knowledge in a database with a similar structure to the target data, such as utilizing Retrieval-Augmented Generation (RAG). RAG combines retrieval systems with generative language models, and it enhances the model’s capabilities by providing access to an external library (Guo et al. 2017; Song et al. 2019; Shi et al. 2018). For dataset retrieval, some existing works (Fleischer et al. 2024; Siriwardhana et al. 2023) use the retrieved documents as contextual information for generation and significantly enrich the model’s knowledge base.
@@
However, this direct method often proves impractical due to the high computational costs of embedding entire domains or whole datasets. Additionally, the retrieved knowledge may not align effectively with the intended application. Such misalignments typically result from variations in data distribution, feature spaces, or contextual differences between the source and target data. Furthermore, the quality of the retrieved data, such as the noise or incomplete features, can also hinder its effectiveness for augmenting data via knowledge augmentation. Thus, in such situations, identifying the optimal knowledge source that provides the most relevant and high-quality data tailored to the specific needs of the target data through the capabilities of LLMs remains a substantial challenge.
@@
Our Targets. Comprehensively, we aim to address three main challenges in retrieving source domain data: (1) how to effectively extract key information based on the target domain with limited computational costs; (2) how to design an efficient automatic source domain retrieval for knowledge augmentation; and (3), how to automatically harmonize the retrieved source data with the target domain to optimize transfer efficiency and improve learning outcomes.
Our Approach. To address the challenges above, we design LLM-Enhanced Knowledge Augmentation (LEKA), a novel and automated data retrieval and augmentation method by knowledge augmentation. Specifically, (1) we utilize an LLM to extract the key textual information of the target domain; (2) we deploy dataset RAG in an external database to efficiently extract relevant knowledge to target domain dataset. The RAG libraries can contain large amounts of continuously updated datasets, ensuring that the data retrieved from these libraries is more precise and timely; (3) then the LLM automatically harmonizes the retrieved source data with the target domain on feature space and marginal probability measures to enhance downstream task learning performance. The data harmonization reduces the structural and semantic differences between the source and target data. This data harmonization is crucial for reducing domain shift, which optimizes transfer efficiency and improves the overall learning outcomes of the model.
As shown in Figure 1, for a target domain targeting predicting a rare cancer, direct learning is challenging due to insufficient sample information. To augment the data with external information, we need human experts with knowledge of both cancer physiological indicators and machine learning to retrieve datasets. However, our method can automate the retrieval of a common cancer dataset as the knowledge source with extensive feature information. This retrieved dataset has a more complete and transparent feature space with a structure similar to the target domain. The LLM then adjusts the retrieved data based on the characteristics of the target domain. In this way, we achieve a highly harmonized source dataset for knowledge augmentation.
@@
Related Work
Knowledge Transfer
Knowledge transfer is a knowledge augmentation method that improves learning on new tasks by transferring knowledge from a related task (Alyafeai, AlShaibani, and Ahmad 2020; Iman, Arabnia, and Rasheed 2023; Bu et al. 2021). Knowledge transfer allows cross-domain knowledge transformation despite data distribution or feature space differences. Specifically, knowledge transfer adapts models developed for one task to perform better on different but related tasks, as it adjusts the feature mappings and decision boundaries to suit new knowledge (He et al. 2021; Han, Pang, and Wu 2021; Yordanov et al. 2021). Finding suitable pre-training knowledge is crucial for knowledge transfer because it significantly enhances the effectiveness of the model by providing a relevant starting point. While many methods exist to adjust existing knowledge within the same or across different domains to improve transfer learning outcomes, the process of retrieving certain knowledge still heavily relies on manual effort.
Data Harmonizing with LLMs
Data harmonization using LLMs offers significant benefits by leveraging their natural language capabilities to standardize and integrate diverse datasets and further enhance model performance through improved data consistency (Cao, Gorodetsky, and Mitkas 2009; Liu et al. 2021; Durante et al. 2024; Zhang et al. 2024d). However, this approach has several challenges, including the high computational costs of training LLMs and potential biases in training data, which can adversely affect the harmonization process (Feng et al. 2021; Li, Hou, and Che 2022; Bu et al. 2019). Furthermore, the risk of overfitting remains a concern, as models may become overly specialized in training data nuances, reducing their effectiveness on new datasets.
Retrieval Augmented Generation for Knowledge Augmentation
Retrieval Augmented Generation (RAG) (Lewis et al. 2020) for knowledge augmentation is a method that enhances the capabilities of generative language models by integrating information retrieval with model generation (Hu and Lu 2024). This technique aids LLMs in tasks that demand deep and specific domain knowledge (Zhang et al. 2024b; Huang and Huang 2024; Zhang et al. 2024a). RAG for knowledge augmentation provides access to expansive external libraries (collections of documents or domain-relevant knowledge), making it especially effective for transferring knowledge across different domains (Siriwardhana et al. 2023; Dan, Lu, and Li 2024). RAG for knowledge augmentation embeds extensive databases directly into the generative process so that specific, domain-related information is both accessible and effectively utilized. In this way it can significantly enhance the performance for complex tasks in knowledge augmentation scenarios.
Preliminary
Definition
@@
Source Domain. The source domain $\mathcal{D}_{S}$ consists of instances paired with labels $y$, represented as $\mathcal{D}_{S}=\{(\mathbf{x},y)|\mathbf{x}_{i}\in\mathcal{X}_{S},y_{i}\in\mathcal{Y}_{S},i=1,\ldots,n^{S}\}$. Here, $\mathcal{X}_{S}$ is the feature space and $\mathcal{Y}_{S}$ is the label space for the source domain. This domain provides labeled data used to train models in preparation for transfer learning tasks.
Target Domain. The target domain $\mathcal{D}_{T}$ typically contains a mix of unlabeled instances and a smaller set of labeled instances, denoted as $\mathcal{D}_{T}=\{\mathbf{x}\in\mathcal{X}_{T}\}\cup\{(\mathbf{x},y)|(\mathbf{x},y)\in\mathcal{X}_{T}\times\mathcal{Y}_{T}\}$. Here, $\mathcal{X}_{T}$ is the feature space and $\mathcal{Y}_{T}$ is the label space for the target domain. We aim to evaluate and fine-tune the transfer learning models with data in the target domain.
@@
Methodology
Figure 2: Framework of LEKA includes: 1) an LLM extracts and embeds the textual information of target dataset, then 2) retrieves datasets in libraries and 3) process data harmonization. With harmonized dataset, we can transfer knowledge from the source dataset we construct to enhance target dataset learning.
@@
Dataset Retrieval
In this phase, our primary objective is to identify and retrieve a source dataset $\mathcal{D}_{S}$ similar to the target dataset $\mathcal{D}_{T}$ for effective pre-training in knowledge transfer scenarios. To achieve this, we utilize an LLM $\mathcal{A}$ to analyze and retrieve data that matches the structural properties and purposes of $\mathcal{D}_{T}$. A straightforward method is to embed the target dataset $\mathcal{D}_{T}$ by $\mathcal{A}$ to capture its essential features. However, this approach can be computationally intensive and prone to errors. Inspired by the strategies outlined in (Zhang et al. 2024c), we turn to leverage the textual information, including the data descriptions $\mathcal{C}_{T}$ and feature names $l_{T}$ of $\mathcal{D}_{T}$. These textual elements provide rich insights into the structure and purpose of $\mathcal{D}_{T}$, and the LLM can precisely focuses on the semantic content of $\mathcal{D}_{T}$.
@@
We then retrieve through in a library $\mathcal{L}$ of dataset for the top-$k$ most relevant datasets $\{\mathcal{D}_{k}\}$ to query $\mathcal{Q}_{t}$. We evaluate the relevance of all datasets in $\mathcal{L}$ to the query $\mathcal{Q}$ by the cosine similarity of the embeddings and select the top-$k$ datasets with the highest similarity:
@@
Data Harmonization
Now we have $k$ potential source domain datasets, we turn to align these datasets with the target dataset $\mathcal{D}_{T}$ with an LLM. This alignment process is to adjust the feature space and marginal probability measures of the source dataset $\mathcal{D}_{k}$ to closely match those of the target dataset $\mathcal{D}_{T}$. This process involves transformations of features and adjustments of distributions. Here we take $\mathcal{D}_{k}$ as an example.
@@
We denote $\mathcal{X}_{k}$ and $\mathcal{X}_{T}$ to be the feature spaces of the source dataset $\mathcal{D}_{k}$ and the target dataset $\mathcal{D}_{T}$, respectively. Our LLM constructs a mapping function $f_{map}:\mathcal{X}_{k}\rightarrow\mathcal{X}_{T}$ for alignment. This function transforms the features in $\mathcal{X}_{k}$ to minimize the distance between the transformed source features and the target features. Here, we aim to minimize the distance $d(f(\mathcal{X}_{k}),\mathcal{X}_{T})$, where $d$ is a kernel distance:
@@
Our LLM then aligns the marginal probability measures $\mathcal{P}_{k}$ and $\mathcal{P}_{T}$ of the source and target datasets. The LLM analyzes and refines the textual information including labels, feature names, and classification probability distribution,.etc. To measure the differences between $\mathcal{D}_{k}$ and $\mathcal{D}_{T}$, we adopt the Wasserstein distance:
@@
With this source-target alignment process, the LLM improves the efficiency of knowledge transfer from the source domain to the target domain. Thus it enhances the overall performance of our transfer learning models. Further, we optimize data handling and boost the adaptability and accuracy of transfer learning to be adopted in various applications.
Knowledge Transfer and Evaluation
@@
During update, the optimization of model parameters $\theta$ leverages both the target dataset $\mathcal{D}_{T}$ and the aligned source dataset $\mathcal{D}_{k}^{\prime}$, which incorporates domain-specific characteristics to reduce domain shift. We define the optimization process as:
@@
This formula aims to fine-tune the decision function $f_{T_{j}}$ by minimizing the weighted sum of expected losses across the datasets. The loss function $\mathcal{L}(f_{T_{j}}(x;\theta),y)$ evaluates prediction accuracy, guiding the adjustment of parameters. The expectations $\mathbb{E}_{(x,y)\in\mathcal{D}_{k}^{\prime}}$ and $\mathbb{E}_{(x,y)\in\mathcal{D}_{T}}$ represent the mean losses over the source and target datasets, respectively. The weighting factor $\alpha$ adjusts the relative influence of each dataset, and keeps flexible adaptation between leveraging established knowledge from the source and integrating new data from the target. In this way, we utilize the transferred knowledge to enhance model performance on the target task.
@@
Experiments
In this section, we present four experiments to demonstrate the effectiveness and impacts of the LEKA. First, we compare the performance of the LEKA against several baseline methods on four downstream tasks. Second, we present the correlations between several target domains and their retrieved source domain. Finally, we discuss the reason for the improvement of performance.
Table 2: Performance comparison of transfer learning methods.

### v2

@@ Abstract
Humans excel in analogical learning and knowledge transfer and, more importantly, possess a unique understanding of identifying appropriate sources of knowledge. From a model’s perspective, this presents an interesting challenge. If models could autonomously retrieve knowledge useful for transfer or decision-making to solve problems, they would transition from passively acquiring to actively accessing and learning from knowledge. However, filling models with knowledge is relatively straightforward—it simply requires more training and accessible knowledge bases. The more complex task is teaching models about which knowledge can be analogized and transferred. Therefore, we design a knowledge augmentation method, LEKA, for knowledge transfer that actively searches for suitable knowledge sources that can enrich the target domain’s knowledge. This LEKA method extracts key information from the target domain’s textual information, retrieves pertinent data from external data libraries, and harmonizes retrieved data with the target domain data in feature space and marginal probability measures. We validate the effectiveness of our approach through extensive experiments across various domains and demonstrate significant improvements over traditional methods in reducing computational costs, automating data alignment, and optimizing transfer learning outcomes.
1 Introduction
@@
Humans are good at identifying proper sources of knowledge. This is an ability that is rooted in our capability for analogical reasoning and knowledge management. In contrast, artificial intelligence models do not inherently possess this intuition: they require explicit instructions and systematic training to identify and utilize relevant information. This gap presents a challenge in enriching domain knowledge and enhancing data augmentation.
@@
Knowledge augmentation is vital for improving model performance, especially in domains with limited information or complex data structures Tang et al. (2020). In knowledge augmentation, knowledge transfer is an important method improving learning performance by transferring knowledge from external information sources Khodaee et al. (2024). In data-limited scenarios, effective knowledge sourcing can bridge domain information gaps and enhance model robustness using relevant external information. In these scenarios, knowledge augmentation, especially knowledge transfer, can reduce reliance on extensive target domain data by strategically selecting sources matching target needs.
@@
Despite its potential, traditional knowledge augmentation methods often involve manual intervention Ringwald and Stiefelhagen (2021); Nam et al. (2024); Zhang et al. (2024d), where human experts select source domains based on their subjective interpretation of the target domain’s requirements. Using tabular learning as an example highlights several challenges: (1) structural and format differences across domains hinder data alignment and integration; (2) discrepancies in tasks and content between the selected source and target domains can diminish the success of knowledge augmentation; and (3) extensive data preprocessing is often required to properly match the chosen source domain dataset. Additionally, relying on human expertise for source domain selection can lead to biases and inefficiencies, as these choices are typically made based on prior knowledge rather than a rigorous, data-driven analysis.
@@
A natural idea is to construct an automated search for datasets with relevant knowledge in a database that has a similar structure to the target data, such as utilizing Retrieval-Augmented Generation (RAG) Gao et al. (2023); Lewis et al. (2020). RAG combines retrieval systems with generative language models, and it enhances the model’s capabilities by providing access to an external library Guo et al. (2017); Zhang et al. (2025b); Shi et al. (2018). For dataset retrieval, some existing works Fleischer et al. (2024); Siriwardhana et al. (2023) use the retrieved documents as contextual information for generation and significantly enrich the model’s knowledge base.
@@
However, this direct method is often impractical due to the high costs of embedding entire domains or datasetsSeemakhupt et al. (2024); Jin et al. (2024). Additionally, retrieved knowledge may not align effectively with the intended applicationEdge et al. (2024); Li and Ramakrishnan (2025). Such misalignments typically result from variations in data distribution, feature spaces, or contextual differences between source and target data. Furthermore, retrieved data quality, like noise or incomplete features, can also hinder its effectiveness for augmenting data via knowledge augmentation. Thus, in such situations, identifying the optimal knowledge source providing the most relevant, high-quality data tailored to specific target data needs via LLM capabilities remains a substantial challenge.
@@
Our Targets.
We aim to address three main challenges in retrieving source domain data: (1) how to effectively extract key information from the target domain with limited computational cost; (2) how to design an efficient, automatic source domain retrieval for knowledge augmentation; and (3) how to automatically harmonize the retrieved source data with the target domain to optimize transfer efficiency and improve learning outcomes.
Our Approach.
To address the challenges above, we design LLM-Enhanced Knowledge Augmentation (LEKA), a novel and automated data retrieval and augmentation method by knowledge augmentation. Specifically, (1) we utilize an LLM to extract the key textual information of the target domain; (2) we deploy dataset-RAG in an external database to efficiently extract knowledge relevant to the target domain dataset. The RAG libraries can contain large amounts of continuously updated datasets, ensuring that the data retrieved from these libraries is more precise and timely; (3) then the LLM automatically harmonizes the retrieved source data with the target domain on feature space and marginal probability measures to enhance downstream task learning performance. The data harmonization reduces the structural and semantic differences between the source and target data. This data harmonization is crucial for reducing domain shift, which optimizes transfer efficiency and improves the overall learning outcomes of the model.
As shown in Figure 1, consider a target domain focused on predicting a rare cancer, direct learning is challenging due to insufficient sample information. To augment the data with external information, would typically require human experts with knowledge of both cancer physiological indicators and machine learning to retrieve datasets. In contrast, our method can automate the retrieval of a common cancer dataset as the knowledge source with extensive feature information. This retrieved dataset possesses a more complete and transparent feature space with a structure similar to the target domain. The LLM then adjusts the retrieved data based on the characteristics of the target domain. In this way, we achieve a highly harmonized source dataset for knowledge augmentation.
@@
2 Related Work
2.1 Knowledge Transfer
Knowledge transfer is a knowledge augmentation method that improves learning on new tasks by transferring knowledge from a related task Alyafeai et al. (2020); Wang et al. (2022, 2025). Knowledge transfer allows cross-domain knowledge transformation despite data distribution or feature space differences. Specifically, knowledge transfer adapts models developed for one task to perform better on different but related tasks, as it adjusts the feature mappings and decision boundaries to suit new knowledge He et al. (2021); Han et al. (2021); Yordanov et al. (2021). Finding suitable pre-training knowledge is crucial for knowledge transfer because it significantly enhances the model’s effectiveness by providing a relevant starting point. While many methods exist to adjust existing knowledge within the same or across different domains to improve transfer learning outcomes, the process of retrieving certain knowledge still heavily relies on manual effort.
2.2 Data Harmonizing with LLMs
Data harmonization using LLMs offers significant benefits by leveraging their natural language capabilities to standardize and integrate diverse datasets and further enhance model performance through improved data consistency Cao et al. (2009); Zhang et al. (2024a); Liu et al. (2021); Durante et al. (2024). However, this approach has several challenges, including the high computational costs of training LLMs and potential biases in training data, which can adversely affect the harmonization process Feng et al. (2021); Xie et al. (2024); Zhang et al. (2025a). Furthermore, the risk of overfitting remains a concern, as models may become overly specialized in training data nuances, reducing their effectiveness on new datasets.
2.3 Retrieval Augmented Generation for Knowledge Augmentation
Retrieval Augmented Generation (RAG) Lewis et al. (2020) for knowledge augmentation is a method that enhances the capabilities of generative language models by integrating information retrieval with model generation Hu and Lu (2024). This technique aids LLMs in tasks that demand deep and specific domain knowledge Zhang et al. (2024b); Huang and Huang (2024). RAG for knowledge augmentation provides access to expansive external libraries (collections of documents or domain-relevant knowledge), making it especially effective for transferring knowledge across different domains Siriwardhana et al. (2023). RAG for knowledge augmentation embeds extensive databases directly into the generative process so that specific, domain-related information is both accessible and effectively utilized. In this way it can significantly enhance the performance for complex tasks in knowledge augmentation scenarios.
3 Preliminary
3.1 Definition
@@
Source Domain. The source domain $\mathcal{D}_{S}$ consists of instances paired with labels $y$, represented as $\mathcal{D}_{S}=\{(\mathbf{x},\mathbf{y})|\mathbf{x}_{i}\in\mathcal{X}_{S},y_{i}\in\mathcal{Y}_{S},i=1,\ldots,n^{S}\}$. Here, $\mathcal{X}_{S}$ is the feature space and $\mathcal{Y}_{S}$ is the label space for the source domain. This domain provides labeled data used to train models in preparation for transfer learning tasks.
Target Domain. The target domain $\mathcal{D}_{T}$ typically contains a mix of unlabeled instances and a smaller set of labeled instances, denoted as $\mathcal{D}_{T}=\{\mathbf{x}\in\mathcal{X}_{T}\}\cup\{(\mathbf{x},\mathbf{y})|(\mathbf{x},\mathbf{y})\in\mathcal{X}_{T}\times\mathcal{Y}_{T}\}$. Here, $\mathcal{X}_{T}$ is the feature space and $\mathcal{Y}_{T}$ is the label space for the target domain. We aim to evaluate and fine-tune the transfer learning models with data in the target domain.
@@ 3.2 Problem Formulation
4 Methodology
Figure 2: Framework of LEKA includes: 1) an LLM extracts and embeds the textual information of the target dataset, then 2) retrieves datasets in libraries and 3) processes data harmonization. With harmonized datasets, we can transfer knowledge from the source dataset we construct to enhance learning on the target dataset.
@@ 4 Methodology
4.1 Dataset Retrieval
In this phase, our primary objective is to identify and retrieve a source dataset $\mathcal{D}_{S}$ similar to the target dataset $\mathcal{D}_{T}$ for effective pre-training in knowledge transfer scenarios. To achieve this, we utilize an LLM $\mathcal{A}$ to analyze and retrieve data that matches the structural properties and purposes of $\mathcal{D}_{T}$. A straightforward method is to embed the target dataset $\mathcal{D}_{T}$ by $\mathcal{A}$ to capture its essential features. However, this approach can be computationally intensive and prone to errors. Inspired by the strategies outlined in Zhang et al. (2024c), we turn to leverage the textual information, including the data descriptions $\mathcal{C}_{T}$ and feature names $l_{T}$ of $\mathcal{D}_{T}$. These textual elements provide rich insights into the structure and purpose of $\mathcal{D}_{T}$, and the LLM can precisely focuses on the semantic content of $\mathcal{D}_{T}$.
@@ 4.1 Dataset Retrieval
We then retrieve from a library $\mathcal{L}$ of datasets for the top-$k$ most relevant datasets $\{\mathcal{D}_{k}\}$ to query $\mathcal{Q}$. We evaluate the relevance of all datasets in $\mathcal{L}$ to the query $\mathcal{Q}$ by the cosine similarity of the embeddings and select the top-$k$ datasets with the highest similarity:
@@
4.2 Data Harmonization
Now that we have $k$ potential source domain datasets, we turn to align these datasets with the target dataset $\mathcal{D}_{T}$ with an LLM. This alignment process is to adjust the feature space and marginal probability measures of the source dataset $\mathcal{D}_{k}$ to closely match those of the target dataset $\mathcal{D}_{T}$. This process involves transformations of features and adjustments of distributions. Here we take $\mathcal{D}_{k}$ as an example.
@@ 4.2 Data Harmonization
We denote $\mathcal{X}_{k}$ and $\mathcal{X}_{T}$ as the feature spaces of the source dataset $\mathcal{D}_{k}$ and the target dataset $\mathcal{D}_{T}$, respectively. Our LLM constructs a mapping function $f:\mathcal{X}_{k}\rightarrow\mathcal{X}_{T}$ for alignment. This function transforms the features in $\mathcal{X}_{k}$ to minimize the distance between the transformed source features and the target features. Here, we aim to minimize the distance $d(f(\mathcal{X}_{k}),\mathcal{X}_{T})$, where $d$ is a kernel distance:
@@
Our LLM then aligns the marginal probability measures $\mathcal{P}_{k}$ and $\mathcal{P}_{T}$ of the source and target datasets. The LLM analyzes and refines the textual information, including labels, feature names, and classification probability distribution, etc. To measure the differences between $\mathcal{D}_{k}$ and $\mathcal{D}_{T}$, we adopt the Wasserstein distance:
@@
With this source-target alignment process, the LLM improves the efficiency of knowledge transfer from the source domain to the target domain. Thus, it enhances the overall performance of our transfer learning models. Further, we optimize data handling and boost the adaptability and accuracy of transfer learning for adoption in various applications
4.3 Knowledge Transfer and Evaluation
@@
During update, the optimization of model parameters $\theta$ leverages both the target dataset $\mathcal{D}_{T}$ and the aligned source dataset $\mathcal{D}_{k}^{\prime}$, a dataset that incorporates domain-specific characteristics to reduce domain shift. We define the optimization process as:
@@
This formula aims to fine-tune the decision function $f_{T_{j}}$ by minimizing the weighted sum of expected losses across the datasets. The loss function $\mathcal{L}(f_{T_{j}}(x;\theta),y)$ evaluates prediction accuracy, guiding the adjustment of parameters. The expectations $\mathbb{E}_{(x,y)\in\mathcal{D}_{k}^{\prime}}$ and $\mathbb{E}_{(x,y)\in\mathcal{D}_{T}}$ represent the mean losses over the source and target datasets, respectively. The weighting factor $\alpha$ adjusts the relative influence of each dataset and enables flexible adaptation between leveraging established knowledge from the source and integrating new data from the target. In this way, we utilize the transferred knowledge to enhance model performance on the target task.
@@
5 Experiments
In this section, we present four experiments to demonstrate the effectiveness and impacts of the LEKA. First, we compare the performance of the LEKA against several baseline methods on four downstream tasks. Second, we present the correlations between several target domains and their retrieved source domains. Finally, we discuss the reason for the improvement of performance.